# Knowledge Adaptation as Posterior Correction

### v1

@@ 5 Conclusion and Discussion
The ability to quickly adapt is the holy grail of intelligence and we propose posterior correction as a natural mechanism for instill this ability in learning algorithms. The suggestion is that, when adapting, the algorithm should aim to correct its posterior distribution. The posterior could be explicit (for example, for a Bayesian learners) or implicit otherwise (as Khan and Rue (2023) have shown). Nevertheless, the suggestion is that all learners should aim to correct this very same object. The proposal shares this view with those of Zellner (1988) who derived Bayes’ rule as an ‘optimal’ information processing rule. There, the information is neither created nor destroyed, simply preserved. Our proposal extends this principle to approximate posteriors obtained by using the Bayesian learning rule of Khan and Rue (2023). The message fits well into the current trend of deep learning where exact posteriors are impossible to compute. The new proposal makes it possible to apply the existing Bayesian principle to all such new cases.
The paper also discusses posterior-corrections methods. Various approximations with Taylor’s method seem useful when computation and memory are limited, but new primal-dual techniques are also attractive to eventually aim for a near-perfect correction. But, two questions still remain: what should the algorithm remember and what new experiences should it seek? We do not provide immediate answers to these but do set the scene to search for them. Essentially, memory should be chosen to minimize the corrections needed in the future, while new experiences should be chosen to enable easy-enough corrections (or at least not too daunting). Both seem extremely difficult because the future is unknown, but an algorithm still has the freedom to explore the brave new world by choosing its future and fixing its past. Posterior correction sets the stage for all such explorations for algorithms that could some day have a natural human-like ability to adapt.
@@ Acknowledgement
I would like to thank my family for their support throughout this work which started in 2019, and continued through pandemic. This work is supported by JST CREST Grant Number JP-MJCR2112. I benefited immensely by having many insightful discussions with many past and current members of the ABI team at RIKEN AIP as well as the CREST-ANR Bayes-Duality Project. I specifically want to thank my colleagues Dr. T. Möllenhoff (RIKEN AIP), Dr. Siddharth Swaroop (Harvard University), Ang Ming Liang (University College London), Nico Daheim (TU Darmstadt), Dharmesh Tailor (University of Amsterdam). Many thanks to Prof. Håvard Rue (KAUST) for his mentorship and encouragement throughout.
Appendix A Expected Hessian for a Bregman Loss
For Bregman losses, we know that the Hessian involves a Generalized Gauss-Newton (GGN) matrix and the Hessian of the model output (Martens, 2020, Eq. 6). Using this, we can then apply delta method to simplify as follows,
@@ Appendix A Expected Hessian for a Bregman Loss
$\begin{split}\mathbf{H}_{i|t}&=\mathbb{E}_{q_{t}}\left[\nabla^{2}\mathcal{L}[y_{i},\hat{y}_{i}(\boldsymbol{\theta})]\right]\\ &=\mathbb{E}_{q_{t}}\left[\nabla f_{i}(\boldsymbol{\theta})\,\,{\left.\nabla^{2}\mathcal{L}[y_{i},z]\right|_{z=f_{i}(\boldsymbol{\theta})}}\,\,\nabla f_{i}(\boldsymbol{\theta})^{\top}+{\left.\nabla\mathcal{L}[y_{i},z]\right|_{z=f_{i}(\boldsymbol{\theta})}}\,\,\nabla^{2}f_{i}(\boldsymbol{\theta})\right]\\ &\approx\nabla f_{i}(\mathbf{m}_{t})\beta_{i|t}\nabla f_{i}(\mathbf{m}_{t})^{\top}+r_{i|t}\nabla^{2}f_{i}(\mathbf{m}_{t}).\end{split}$ (38)

### v3

@@ 5 Conclusion and Discussion
The ability to quickly adapt is the holy grail of intelligence and we propose posterior correction as a mechanism to instill this ability in learning algorithms. The suggestion is that, when adapting, the algorithm should aim to correct its posterior distribution. The proposal shares this view with those of zellner1988optimal who derived Bayes’ rule as an ‘optimal’ information processing rule. There, the information is neither created nor destroyed, simply preserved. Our proposal extends this principle to approximate posteriors obtained by using the Bayesian learning rule of khanRue23. The goal now is to preserve the information and to maintain optimality within a class of candidate posteriors. The corrections ensure exactly this. The message fits well into the current trend of deep learning where exact posteriors are impossible to compute, but the computation of approximate posterior is still feasible. The new proposal makes it possible to apply the existing Bayesian principle to all such new cases.
The paper also presents methods to correct the posterior. Various approximations with Taylor’s method seem useful when computation and memory are limited, but new primal-dual techniques are also attractive to eventually aim for a near-perfect correction. But, two questions still remain: what should the algorithm remember and what new experiences should it seek? We do not provide immediate answers to these but do set the scene to search for them. Essentially, memory should be chosen to minimize the corrections needed in the future, while new experiences should be chosen to enable easy-enough corrections (or at least not too daunting). Both seem extremely difficult because the future is unknown, but an algorithm still has the freedom to explore the brave new world by choosing its future and fixing its past. Posterior correction sets the stage for all such explorations for algorithms that could some day have a natural human-like ability to adapt.
@@ Acknowledgement
I thank my family for their support throughout this work which started in 2019 and continued through the pandemic. This work is supported by JST CREST Grant Number JP-MJCR2112. I benefited immensely from many insightful discussions with many past and current members of the ABI team at RIKEN AIP as well as the CREST-ANR Bayes-Duality Project. I specifically want to thank my colleagues Dr. T. Möllenhoff (RIKEN AIP), Dr. Siddharth Swaroop (Harvard University), Ang Ming Liang (University College London), Nico Daheim (TU Darmstadt) and Dharmesh Tailor (University of Amsterdam). Many thanks to Prof. Håvard Rue (KAUST) for his mentorship and encouragement throughout.
Appendix A Examples of Dual Parameterization of the VB Posterior
A.1 A One-Dimensional Logistic Regression Example
We give details on the illustration given in Fig. 2 for a binary logistic-regression example. This is a 1-D example with a scalar input $x$ with a label $y\in\{0,1\}$. We have three input-output pairs $(x_{i},y_{i})$, one for the square class (call it class 0) and two for the circle class (call it class 1). This is shown in the first panel of Fig. 5. A classifier is simply a vertical line located at $\theta$. For instance, we can say that when $x>\theta$ then $y=1$, otherwise $y=0$. The logistic loss provides a ‘soft’ thresholding via a sigmoid function,
$\mbox{$\sigma$}(x-\theta)=\frac{1}{1+\exp(-(x-\theta))}.$
This is then used to define the likelihood function,
$p(y_{i}|x_{i},\theta)=\mbox{$\sigma$}(f_{i})^{y_{i}}\left[1-\mbox{$\sigma$}(f_{i})\right]^{1-y_{i}},$
where $f_{i}=x_{i}-\theta$ are example dependent ‘regression’ functions, also known as logits. The loss function is simply the negative-log of the likelihood,
$\ell_{i}(\theta)=-\log p(y_{i}|x_{i},\theta)=-y_{i}\log\mbox{$\sigma$}(f_{i})-(1-y_{i})\left[1-\mbox{$\sigma$}(f_{i})\right].$
This is the two class version of the well-known cross-entropy loss. The likelihood and loss functions are visualized in the second and third panel of Fig. 5 for $y=1$ as a function of $x$.
Figure 5: One dimensional binary classification, and its likelihood and loss function.
The goal of learning is to find $\theta$ such that all examples are classified well, which can be formulated as a (regularizerd) empirical risk minimization,
$\theta_{t}=\arg\min_{\theta}\sum_{i=1}^{t}\ell_{i}+\mbox{$\frac{1}{2}$}\theta^{2}.$
The first panel in Fig. 6 below illustrates all the losses $\ell_{i}$ while the second panel shows the ERM. Note that the $x$-axis here represents $\theta$ in contrast to Fig. 5 where we plotted the likelihood functions against $x$. The second panel shows the overall loss and minimizer which is located between $x_{1}$ and $x_{2}$.
Figure 6: The first panel shows the losses $\ell_{i}$ for the three examples. The second panel shows the overall ERM and the minimizer $\theta_{t}$. The third panel shows a Gaussian posterior $q(\theta)$ with mean $m=\theta_{t}$. The fourth panel shows the expectation of the ERM over $q_{t}$.
The third panel illustrates a Bayesian approach where the goal is to locate a Gaussian posterior $q(\theta)=\mbox{${\cal N}$}(\theta|m,\sigma^{2})$ whose mean is close to $\theta_{t}$. The resulting loss function uses an expected loss $\mathbb{E}_{q}[\ell]$ which can be plotted as a function of the mean $m$. This is shown in the fourth panel where we see that the expected loss closely follows the ERM loss. The higher values are due to averaging over the neighborhood of $m$. The minimizer $m_{t}$ is close to $\theta_{t}$ which can be recovered from the VB objective by essentially ignoring the expectation (via the delta method as explained in the main text).
We will now derive the form of the site functions and show that the Gaussian posterior can be written as a product of the three sites. We denote the $L_{2}$ regularizer $\ell_{0}=\mbox{$\frac{1}{2}$}\theta^{2}$. As shown in Table 1, the site function for a Gaussian $q_{t}=\mbox{${\cal N}$}(m_{t},\sigma_{t}^{2})$ is
$\hat{\ell}_{i|t}=\theta\cdot\mathbb{E}_{q_{t}}[\nabla\ell_{i}]+\mbox{$\frac{1}{2}$}(\theta-m_{t})^{2}\cdot\mathbb{E}_{q_{t}}[\nabla^{2}\ell_{i}]+\text{const.}$
The expectations of the gradient and Hessian of the logistic loss can be simplified by noting the fact that $\mbox{$\sigma$}^{\prime}(f)=\mbox{$\sigma$}(f)[1-\mbox{$\sigma$}(f)]$, as shown below
@@ A.1 A One-Dimensional Logistic Regression Example
$\begin{split}\mathbb{E}_{q_{t}}[\nabla\ell_{i}(\theta)]&=\mathbb{E}_{q_{t}}\left[\left(\frac{y_{i}}{\mbox{$\sigma$}(f_{i})}-\frac{1-y_{i}}{1-\mbox{$\sigma$}(f_{i})}\right)\mbox{$\sigma$}^{\prime}(f_{i})\right]=\left[y_{i}-\mathbb{E}_{q_{t}}\left(\mbox{$\sigma$}(f_{i})\right)\right]:=\alpha_{i|t}\\ \mathbb{E}_{q_{t}}[\nabla^{2}\ell_{i}(\theta)]&=-\mathbb{E}_{q_{t}}\left[\nabla\mbox{$\sigma$}(f_{i})\right]=\mathbb{E}_{q_{t}}\left[\mbox{$\sigma$}^{\prime}(f_{i})\right]:=\beta_{i|t}\end{split}$ (51)
For simplicity, we denote the two quantities by $\alpha_{i|t}$ and $\beta_{i|t}$. We note that $\alpha_{i|t}\in(-1,1)$ is also the residual and $\beta_{i|t}\in(0,1/4)$, a strictly positive scalar.
The site for the loss and likelihood can respectively be written as
$\hat{\ell}_{i|t}=\theta\alpha_{i|t}+\mbox{$\frac{1}{2}$}\beta_{i|t}(\theta-m_{t})^{2}+\text{const.}=\mbox{$\frac{1}{2}$}\beta_{i|t}\left(\theta-\tilde{y}_{i|t}\right)^{2}+\text{const.}$
$\implies\exp(-\hat{\ell}_{i|t})\propto\mbox{${\cal N}$}(\theta|\tilde{y}_{i|t},1/\beta_{i|t})$
where $\tilde{y}_{i|t}=m_{t}-\alpha_{i|t}/\beta_{i|t}$. Fig. 7 illustrates the site function of $x_{1}$ which is located on the left side of $m_{t}$.
Figure 7: The site functions (in dark blue) for the loss (left) and likelihood (right), both shown in light blue.
The pair $(\alpha_{i|t},\beta_{i|t})$ determine the location and curvature, respectively. Essentially, the location of the quadratic shown on the left is $|\frac{\alpha_{i|t}}{\beta_{i|t}}|$ distance away from the mean $m_{t}$. Whether it is on the left or right depends on the sign of the ratio. The variance of the Gaussian factor is equal to $1/\beta_{i|t}$. The quantity $\tilde{y}_{i|t}$ can be seen as the location or as a pseudo-Gaussian observation. Note that the Gaussian factor over $\theta$ is not always guaranteed to be a valid probability ‘distribution’ over $\theta$.
Fig. 8 compares the sites for the losses and likelihoods of all the three example. The site of the observations $x_{1}$ and $x_{2}$ both which are close to $m_{t}$ have a sharper curvature than $x_{3}$ which is far away. The location of the observation also decides the location of the sites. The input $x_{1}$ is on the left side of $m_{t}$ and its site is located on the other side, while the opposite is true for $x_{2}$. The site for $x_{3}$ is located much farther from $m_{t}$.
Figure 8: Each row shows the sites of an example. Sites for the loss are shown on the left while those for the likelihood are on the right.
From the figure above it is clear that the sites of $x_{1}$ and $x_{2}$ are much more prominent than that of $x_{3}$. As a result, the contributions of each site to $q_{t}$ also differs drastically from each other. We will now present precise expressions that show that the example $(x_{3},y_{3})$ contributes a lot less to $q_{t}$ than the other two examples. In fact, the decision would change very little even if $(x_{3},y_{3})$ were to be removed. To see this, let us write the dual representation of the posterior in terms of the site functions,
$\mbox{${\cal N}$}(\theta|m_{t},\sigma_{t}^{2})$ $=\frac{1}{\mathcal{Z}_{t}}\exp\left(-\mbox{$\frac{1}{2}$}\theta^{2}\right)\prod_{i=1}^{t}\exp\left[-\mbox{$\frac{1}{2}$}\beta_{i|t}\theta^{2}+\tilde{y}_{i|t}\beta_{i|t}\theta\right]$ (52)
This is illustrated in the figure below (where we ignored the prior term) to show how Gaussian factors are multiplied together to find the posterior over $\theta$.
Figure 9: Multiplication of the sites followed by normalization yields the posterior.
We will express the mean $m_{t}$ and variance $\sigma_{t}^{2}$ in terms of the site parameters $(\alpha_{i|t},\beta_{i|t})$. Both sides in Eq. 52 have Gaussian forms, therefore we can expand them to see how each of them contributes to the sufficient statisics $(\theta,\theta^{2})$ of the posterior,
$\exp\left[-\mbox{$\frac{1}{2}$}\left(\frac{1}{\sigma_{t}^{2}}\right)\theta^{2}+\left(\frac{m_{t}}{\sigma_{t}^{2}}\right)\theta\right]$ $=\frac{1}{\mathcal{Z}_{t}}\exp\left[-\mbox{$\frac{1}{2}$}\theta^{2}\right]\prod_{i=1}^{t}\exp\left[-\mbox{$\frac{1}{2}$}\beta_{i|t}\theta^{2}+\tilde{y}_{i|t}\beta_{i|t}\theta\right]$
$=\frac{1}{\mathcal{Z}_{t}}\exp\left[-\mbox{$\frac{1}{2}$}\left(1+\sum_{i=1}^{t}\beta_{i|t}\right)\theta^{2}+\left(\sum_{i=1}^{t}\tilde{y}_{i|t}\beta_{i|t}\right)\theta\right]$
$=\frac{1}{\mathcal{Z}_{t}}\exp\left[-\mbox{$\frac{1}{2}$}\left(1+\sum_{i=1}^{t}\beta_{i|t}\right)\theta^{2}+\left(m_{t}\sum_{i=1}^{t}\beta_{i|t}-\sum_{i=1}^{t}\alpha_{i|t}\right)\theta\right].$
Matching the terms in front of $\theta^{2}$, we find that the posterior precision $1/\sigma_{t}^{2}$ is expressed as the sum of individual precisions $\beta_{i|t}$, that is,
$\frac{1}{\sigma_{t}^{2}}=1+\sum_{i=1}^{t}\beta_{i|t}.$ (53)
Then matching the terms in front of $\theta$ and using the last expression, we get
$\frac{m_{t}}{\sigma_{t}^{2}}$ $=m_{t}\sum_{i=1}^{t}\beta_{i|t}-\sum_{i=1}^{t}\alpha_{i|t}=m_{t}\left(\frac{1}{\sigma_{t}^{2}}-1\right)-\sum_{i=1}^{t}\alpha_{i|t}\quad$ $\implies\quad m_{t}=-\sum_{i=1}^{t}\alpha_{i|t}$ (54)
Readers familiar with the representer theorem may see the above formula as an instance of the well-known representer theorem from Kernel methods (scholkopf2001generalized). This is now recovered by matching the sufficient statistics of the posterior with those of the dual form. The dual form presented here is more general because it also yields a representation formula for the variance as shown in Eq. 53.
A.2 Binary Logistic Regression: General Case
Consider a binary logistic regression problem with $N$ input-output pairs $(\mbox{$\mbox{$\mathbf{x}$}$}_{i},y_{i})$. For this case, we can define the logits as linear predictors $f_{i}=\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}$. We will derive the form of the site functions for a full Gaussian distribution $q_{t}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{t},\mathbf{\Sigma}_{t})$. The site function are defined through $\mathbb{E}_{q_{t}}[\nabla\ell_{i}]$ and $\mathbb{E}_{q_{t}}[\nabla^{2}\ell_{i}]$, both of which can be simplified. For example, we can change the variable from $\boldsymbol{\theta}$ to $f$ to write
$\mathbb{E}_{q_{t}}[\nabla\ell_{i}]=\int\underbrace{\mbox{$\mbox{$\mathbf{x}$}$}_{i}\ell_{i}^{\prime}(\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta})}_{=\nabla\ell_{i}(\boldsymbol{\theta})}\underbrace{\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{t},\mathbf{\Sigma}_{t})}_{=q_{t}(\boldsymbol{\theta})}d\boldsymbol{\theta}=\mbox{$\mbox{$\mathbf{x}$}$}_{i}\int\ell_{i}^{\prime}(f)\underbrace{\mbox{${\cal N}$}(f|\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\mathbf{m}_{t},\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\mathbf{\Sigma}_{t}\mbox{$\mbox{$\mathbf{x}$}$}_{i})}_{=\tilde{q}_{i|t}(f)}df=\mbox{$\mbox{$\mathbf{x}$}$}_{i}\bar{\alpha}_{i|t}.$
Here, $\tilde{q}_{i|t}(f)$ is a distribution over $f=\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}$ derived from $q_{t}(\boldsymbol{\theta})$ and we define
$\bar{\alpha}_{i|t}=\mathbb{E}_{\tilde{q}_{i|t}}[\ell_{i}^{\prime}]\text{ where }\ell_{i}^{\prime}(f_{i})=p(f_{i})-y_{i}$
is the residual. In a similar way, we can write $\mathbb{E}_{q_{t}}[\nabla^{2}\ell_{i}]=\mbox{$\mbox{$\mathbf{x}$}$}_{i}\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\bar{\beta}_{i|t}$, where we define
$\bar{\beta}_{i|t}=\mathbb{E}_{\tilde{q}_{i|t}}[\ell_{i}^{\prime\prime}]\text{ with }\ell_{i}^{\prime\prime}(f_{i})=p(f_{i})[1-p(f_{i})].$
Thus, we get write the site function as 1-dimension function over $f_{i}(\boldsymbol{\theta})=\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}$,
$\begin{split}\hat{\ell}^{\text{full}}_{i|t}(\boldsymbol{\theta})&=\boldsymbol{\theta}^{\top}\mbox{$\mbox{$\mathbf{x}$}$}_{i}\bar{\alpha}_{i|t}+\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\mbox{$\mbox{$\mathbf{x}$}$}_{i}\bar{\beta}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}(\boldsymbol{\theta}-\mathbf{m}_{t})\\ &=\bar{\alpha}_{i|t}f_{i}(\boldsymbol{\theta})+\mbox{$\frac{1}{2}$}\bar{\beta}_{i|t}(f_{i}(\boldsymbol{\theta})-\widetilde{m}_{i|t})^{2}+\text{const.}\\ &=\mbox{$\frac{1}{2}$}\beta_{i|t}(f_{i}(\boldsymbol{\theta})-\tilde{y}_{i|t})^{2}+\text{const.}\end{split}$ (55)
Here, we defined $\widetilde{m}_{i|t}=\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\mathbf{m}_{t}$ and $\tilde{y}_{i|t}=\widetilde{m}_{i|t}-\bar{\alpha}_{i|t}/\bar{\beta}_{i|t}$. Similarly to the simple 1-dimensional case, the site is a quadratic function whose location is $|\bar{\alpha}_{i|t}/\bar{\beta}_{i|t}|$ distance away from $\widetilde{m}_{i|t}$ and its curvature is $\bar{\beta}_{i|t}$. The site can also be written as a Gaussian factor,
$\exp(-\hat{\ell}_{i|t})\propto\mbox{${\cal N}$}(\theta|\tilde{y}_{i|t},1/\bar{\beta}_{i|t})$
With this site function, we can write the dual form of the posterior by using Eq. 3 (and assuming a standard normal prior),
$\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{t},\mathbf{\Sigma}_{t})$ $=\frac{1}{\mathcal{Z}_{t}}\exp\left(-\mbox{$\frac{1}{2}$}\boldsymbol{\theta}^{\top}\boldsymbol{\theta}\right)\prod_{i=1}^{t}\exp\left[-\mbox{$\frac{1}{2}$}\bar{\beta}_{i|t}f_{i}(\boldsymbol{\theta})^{2}+\tilde{y}_{i|t}\bar{\beta}_{i|t}f_{i}(\boldsymbol{\theta})\right]$
We can also express the mean $\mathbf{m}_{t}$ and covariance $\mathbf{\Sigma}_{t}$ in terms of the site parameters $(\bar{\alpha}_{i|t},\bar{\beta}_{i|t})$. There are two ways to go about it by using the two forms given in Eq. 36, one using the natural parameter form and the other using the distributional form. Here, we will use the latter and match terms in front of the sufficient statistics $\boldsymbol{\theta}$ and $\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$, as shown below:
$\exp\left[-\mbox{$\frac{1}{2}$}\boldsymbol{\theta}^{\top}\mathbf{\Sigma}_{t}^{-1}\boldsymbol{\theta}+\mathbf{m}_{t}^{\top}\mathbf{\Sigma}_{t}^{-1}\boldsymbol{\theta}\right]$ $=\frac{1}{\mathcal{Z}_{t}}\exp\left[-\mbox{$\frac{1}{2}$}\boldsymbol{\theta}^{\top}\mbox{$\mbox{$\mathbf{I}$}$}\boldsymbol{\theta}\right]\prod_{i=1}^{t}\exp\left[-\mbox{$\frac{1}{2}$}\boldsymbol{\theta}^{\top}\mbox{$\mbox{$\mathbf{x}$}$}_{i}\bar{\beta}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}+\tilde{y}_{i|t}\bar{\beta}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}\right]$
$=\frac{1}{\mathcal{Z}_{t}}\exp\left[-\mbox{$\frac{1}{2}$}\boldsymbol{\theta}^{\top}\left(\mbox{$\mbox{$\mathbf{I}$}$}+\sum_{i=1}^{t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}\bar{\beta}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\right)\boldsymbol{\theta}+\left(\sum_{i=1}^{t}\tilde{y}_{i|t}\bar{\beta}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}\right)\boldsymbol{\theta}\right]$
Matching the coefficients of $\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$ gives us,
$\mathbf{\Sigma}_{t}^{-1}=\mbox{$\mbox{$\mathbf{I}$}$}+\sum_{i=1}^{t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}\bar{\beta}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}.$ (56)
Similarly, matching the coefficients of $\boldsymbol{\theta}$, we get
$\mathbf{\Sigma}_{t}^{-1}\mathbf{m}_{t}=\sum_{i=1}^{t}(\bar{\beta}_{i|t}\widetilde{m}_{i|t}-\bar{\alpha}_{i|t})\mbox{$\mbox{$\mathbf{x}$}$}_{i}=\left(\sum_{i=1}^{t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}\bar{\beta}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\right)\mathbf{m}_{t}-\sum_{i=1}^{t}\bar{\alpha}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}=\left(\mathbf{\Sigma}_{t}^{-1}-\mbox{$\mbox{$\mathbf{I}$}$}\right)\mathbf{m}_{t}-\sum_{i=1}^{t}\bar{\alpha}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}$
This gives us the following representer theorem,
$\mathbf{m}_{t}=-\sum_{i=1}^{t}\bar{\alpha}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}.$ (57)
This result is analogues to the well-known representer theorem, specialized to finite dimensional features. The form used there can be obtained by simply applying an arbitary test input $\mbox{$\mbox{$\mathbf{x}$}$}_{*}^{\top}$ on both sides, and defining $f_{*|t}=\mbox{$\mbox{$\mathbf{x}$}$}_{*}^{\top}\mathbf{m}_{t}$ and $\kappa(\mbox{$\mbox{$\mathbf{x}$}$}_{*},\mbox{$\mbox{$\mathbf{x}$}$}_{i})=\mbox{$\mbox{$\mathbf{x}$}$}_{*}^{\top}\mbox{$\mbox{$\mathbf{x}$}$}_{i}$. This gives us
$f_{*|t}=-\sum_{i=1}^{t}\bar{\alpha}_{i|t}\kappa(\mbox{$\mbox{$\mathbf{x}$}$}_{*},\mbox{$\mbox{$\mathbf{x}$}$}_{i}).$ (58)
By using delta method, we can approximate $\bar{\alpha}_{i|t}\approx\ell^{\prime}(\widetilde{m}_{i|t})$ which recovers exactly the result of the representer theorem. Here too, our dual form is more general because we also get a representation formula for the covariance, as shown in Eq. 56.
A.3 Dual Forms During Training for Logistic Regression
The dual form given in Eq. 3 also can be written during the iteration of the BLR as shown in Eq. 36. We will now give an example for the logistic regression case.
$q_{t}\leftarrow\frac{1}{\mathcal{Z}_{t}}q_{t}^{1-\rho}\prod_{i=0}^{t}\exp\left(-\rho\hat{\ell}_{i|t}\right)$
As shown in (khanRue23, Sec. 3.1.2), the update can be written in a Newton-like form, by first defining the precondition matrix via the covariance,
$\left(\mathbf{\Sigma}_{t}^{\text{new}}\right)^{-1}\leftarrow(1-\rho)\mathbf{\Sigma}_{t}^{-1}+\rho\left(\mbox{$\mbox{$\mathbf{I}$}$}+\sum_{i=1}^{t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}\bar{\beta}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\right),$
and then using it to update of the mean,
$\mathbf{m}_{t}\leftarrow\mathbf{m}_{t}-\rho\left(\mathbf{\Sigma}_{t}^{\text{new}}\right)^{-1}\left(\sum_{i=1}^{t}\bar{\alpha}_{i|t}\mbox{$\mbox{$\mathbf{x}$}$}_{i}\right)$ (59)
Multiplying both sides by $\mbox{$\mbox{$\mathbf{x}$}$}_{*}^{\top}$, we can write the update in the function space in terms of individual contributions from each data example,
$f_{*|t}\leftarrow f_{*|t}-\rho\sum_{i=1}^{t}\bar{\alpha}_{i|t}\kappa(\mbox{$\mbox{$\mathbf{x}$}$}_{*},\mbox{$\mbox{$\mathbf{x}$}$}_{i})\qquad\text{ where }\quad\kappa(\mbox{$\mbox{$\mathbf{x}$}$}_{*},\mbox{$\mbox{$\mathbf{x}$}$}_{i})=\mbox{$\mbox{$\mathbf{x}$}$}_{*}^{\top}\left(\mathbf{\Sigma}_{t}^{\text{new}}\right)^{-1}\mbox{$\mbox{$\mathbf{x}$}$}_{i}.$
A derivation for neural network is in khan2019approximate by using Gauss-Newton approximation, and a slightly more general form is given in nickl2023memory.
Appendix B Extension of the Dual Form to EF Distributions with Non-Constant Base Measures
For distributions with non-constant base measures $h(\boldsymbol{\theta})$, only a small change in the second step is required. Consider the family $q(\boldsymbol{\theta})=h(\boldsymbol{\theta})\exp(\langle\mathbf{T}(\boldsymbol{\theta}),\boldsymbol{\lambda}\rangle-A(\boldsymbol{\lambda}))$ with the base measure, then the natural gradient of the entropy term (khanRue23, App. B) also contains an additional term $\widetilde{\nabla}\mathbb{E}_{q_{t}}[\log h]$. This can be added in the existing dual form by simply redefining the regularizer site as follows:
$\hat{\ell}_{0|t}(\boldsymbol{\theta})=\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{0}+\log h]\rangle-\log h(\boldsymbol{\theta}).$ (60)
Appendix C Expected Hessian for a Bregman Loss
For Bregman losses, we know that the Hessian involves a Generalized Gauss-Newton (GGN) matrix and the Hessian of the model output (martens2020, Eq. 6). Using this, we can then apply delta method to simplify as follows,
$\begin{split}\bar{\mathbf{H}}_{i|t}&=\mathbb{E}_{q_{t}}\left[\nabla^{2}\mathcal{L}[y_{i},\hat{y}_{i}(\boldsymbol{\theta})]\right]\\ &=\mathbb{E}_{q_{t}}\left[\nabla f_{i}(\boldsymbol{\theta})\,\,{\left.\nabla^{2}\mathcal{L}[y_{i},z]\right|_{z=f_{i}(\boldsymbol{\theta})}}\,\,\nabla f_{i}(\boldsymbol{\theta})^{\top}+{\left.\nabla\mathcal{L}[y_{i},z]\right|_{z=f_{i}(\boldsymbol{\theta})}}\,\,\nabla^{2}f_{i}(\boldsymbol{\theta})\right]\\ &\approx\nabla f_{i}(\mathbf{m}_{t})\beta_{i|t}\nabla f_{i}(\mathbf{m}_{t})^{\top}+r_{i|t}\nabla^{2}f_{i}(\mathbf{m}_{t}).\end{split}$ (61)